OpenStudy (anonymous):

Here's a fun question.\[S_n=\sum_{k=1}^n\left(\frac kn\right)^n\\\lim_{n\to\infty}S_n=\,\,?\]Note: This is a challenge. I have the answer/method if you want to see it.

OpenStudy (lgbasallote):

you remind me of @FoolForMath :C

OpenStudy (anonymous):

I was trying to fill in his absence. :c

OpenStudy (lgbasallote):

it still feels different...his felt like you get humiliated because a "fool" is asking hard questions in yours i feel skeptic...for obvious reasons

OpenStudy (anonymous):

Ahaha.

OpenStudy (anonymous):

I'd like to note that this question is "answerable" using techniques learned in basic calculus, but a more rigorous treatment can be provided using analysis.

OpenStudy (anonymous):

1.58197

OpenStudy (anonymous):

You can express the answer in terms of natural numbers and \(e\).

OpenStudy (anonymous):

But, yes, you are correct. And apologies for multiple posting.

OpenStudy (anonymous):

@KingGeorge You would be interest?

OpenStudy (zarkon):

\[\frac{e}{e-1}\]

OpenStudy (zarkon):

Equivalent into \[\sum_{k=0}^{\infty}e^{-k}\]

OpenStudy (anonymous):

That's right. Do you have a proof written down, or would you like to see it?

OpenStudy (zarkon):

I think the key is to notice that\[\lim_{n\to\infty}\left(\frac{n-m}{n}\right)^n=e^{-m}\]

OpenStudy (anonymous):

Well, it doesn't matter. I'll post the proof: http://mathdl.maa.org/images/cms_upload/Holland-MMz-201039490.pdf

OpenStudy (anonymous):

\[ \frac{e}{e-1.}\approx 1.5819 \]